OpenStudy (anonymous):

Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7

OpenStudy (stormfire1):

Well, the simplest way is to note that 4 6/7 + 1/7 = 5 Then you can see that 9 1/4 + 3/4 = 10 5 + 10 = 15

OpenStudy (saifoo.khan):

im getting 17

OpenStudy (anonymous):

yes thank you((:

OpenStudy (stormfire1):

Ah...I missed the 2 there. Let's try it again... 4 6/7 + 1/7 = 5 9 1/4 + 2 3/4 = 12 So you end up with 12 + 5 = 17